3x+4y=1
2x+3y-z=1
3x+4y-3z=4

Answer by FrankM(1040)

3x+4y+0z=1 Eq1 (i added the 0z to just see the whole line)
2x+3y-z=1 Eq2
3x+4y-3z=4 Eq3
-6x-9y+3z=-3 Eq2 X -3
3x+4y-3z=4 Eq3
-3x-5y+0z=1 sum
3x+4y+0z=1 Eq 1
0x-1y+0z=2 sum
y=-2
3x+4(-2)+0z=1 eq1
3x-8=1
3x=9
x=3
2(3)+3(-2)-z=1 eq 2
6-6+z=1
z=1
3,-2,1